• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 11:01:06 on January29,2002

    Entry number 81225

    keyword=day shift summary

    8:15 Beam is back, continuing with
    4F, LH2, Rad, MON, NoVeto, 35uA
    Runs: 1846

    8:20 beam off
    1847 is cosmic run

    9:15 Beam back to hall.
    Beam quality quite bad from 9:15 to about 9:45. Frequent beam trips.

    10:50 Beam down. Vacuum problem. Shift changed to standby.